ON TWO-GENERATOR SATELLITE KNOTS

STEVEN A. BLEILER AMELIA C. JONES

Abstract. Techniques are introduced which determine the geometric structure of non-simple two-generator 3-manifolds from purely algebraic data. As an application, the satellite knots in the 3-sphere with a two-generator presentation in which at least one generator is represented by a meridian for the knot are classified.